OpenStudy (anonymous):

15x+2x can be written as ______ x(15+2x),(15+2)x,(15x+2x)x

OpenStudy (texaschic101):

(15 + 2)x or x(15 + 2)

OpenStudy (the_fizicx99):

Not technically if it's being raised to a power it completely changes the result, a lot of the times people don't include the " ^ ", (15+2)^x = (17)^x = 17^x Whereas, x(15+2) = 15x + 2x = 17x There's a difference between \(\ 17x \) and \(\ 17^x \)